OpenStudy (anonymous):

Factor the polynomial completely. That is, once you are finished factoring, none of the factors you obtain should be factorable: 6x^2 − 29x + 20

OpenStudy (anonymous):

Lol maan come on i know what you doing. This will be the last one ;)

OpenStudy (anonymous):

6x^(2)-29x+20 For a polynomial of the form ax^(2)+bx+c, find two factors of a*c (120) that add up to b (-29).In this problem -(5)/(6)*-4=(20)/(6) (which is (c)/(a)) and -(5)/(6)-4=-(29)/(6) (which is ((b)/(a)) , so insert -(5)/(6) as the right hand term of one factor and -4 as the right-hand term of the other factor. (x-(5)/(6))(x-4) Remove the fraction by multiplying the first term of the factor by the denominator of the second term. (6x-5)(x-4)

OpenStudy (anonymous):

lmao, it's a online quiz lol I need to properly pass, i'll put in some work on the next one :)